OpenStudy (anonymous):

A box of nutmeg weighs 1.56 grams. If there are 22 teaspoons of nutmeg in the box, how much does one teaspoon of nutmeg weigh? (Round your answer to the nearest thousandth.)

OpenStudy (anonymous):

22 teaspoons weighs 1.56 grams then 1 teaspoon weighs x grams...proportion time... 22/1.56 = 1/x (cross multiply) (22)(x) = (1.56)(1) 22x = 1.56 (divide by 22 on both sides to get x by itself) 22x/22 = 1.56/22 x = .071 ANSWER : One teaspoon of nutmeg weighs .071 grams.

OpenStudy (anonymous):